# Robustness of contextuality under different types of noise as quantifiers for parity-oblivious multiplexing tasks

### v1

@@ II.1 Robustness to depolarising noise in $n$-to-1 parity-oblivious multiplexing
$p(b=x_{y}|P_{x},M_{y})$ $=$ $\operatorname{Tr}(E_{b=x_{y}}\rho_{x})$
$\mapsto$ $(1-r)\operatorname{Tr}(E_{b=x_{y}}\rho_{x})+r\operatorname{Tr}(E_{b=x_{y}}\mu)$
$=$ $(1-r)p(b=x_{y}|P_{x},M_{y})+r\operatorname{Tr}(E_{b=x_{y}}\mu),$
where $E_{b=x_{y}}$ are the POVM elements of measurement $M_{y}$. Once we calculate the success rate for the depolarised version of the experiment, we find
$s_{depol}$ $=$ $\frac{(1-r)}{2^{n}n}\sum_{y=1}^{n}\sum_{x\in[0,1]^{\times n}}p(b=x_{y}|P_{x},M_{y})$ (8)
$\hskip 39.83368pt+\ \frac{r}{2^{n}n}\sum_{y=1}^{n}\sum_{x\in[0,1]^{\times n}}\operatorname{Tr}(E_{b=x_{y}}\mu)$
@@
where the second equality follows from the fact that for every choice of measurement $y$, there will be $2^{n-1}$ strings $x$ with $y$-th input equal to 00, and $2^{n-1}$ strings with the same input equal to $1$. Since all $E_{b=x_{y}}$ are elements of POVM, it must be that $E_{b=0_{y}}+E_{b=1_{y}}=\mathbbm{1}$ for any $y$, simplifying the summation.
@@
With this relation, one can derive bounds on the number of bits that can be optimally encoded in a particular quantum system by assuming bounds on the maximal robustness to depolarising achievable for this system. For instance, for qubits, it is common to assume that $r_{min}^{depol}$ can never exceed the value $\frac{1}{2}$ for whatever signature of quantum advantage taken into consideration, in particular $r_{min}^{depol}<\frac{1}{2}$ for a finite amount of extremal states and effects in the fragment (which is always the case in $n$-to11-1- 1 POM tasks with finite $n$). We provide a compelling numerical induction for this bound in Appendix C.1 that relies on the simplex embeddability of an ever-increasing set of preparations and measurements. Then notice that the optimal quantum success rate is given by $s=\frac{1}{2}(1+\frac{1}{\sqrt{n}})$ [14], while $s_{NC}=\frac{1}{2}(1+\frac{1}{n})$ [4]. If $r_{min}^{depol}<\frac{1}{2}$, then
@@
This result has been derived before by geometrical [47] and locality [48] arguments, but Proposition 1 allows us to reframe it solely on simplicial embedding assumptions. This method can, in principle, be generalized: if one finds bounds on robustness for quantum systems of higher dimensions, one can immediately use Eq. (11) to derive an upper bound to the number of optimally encoded bits.
@@ II.2 Robustness to dephasing noise in 3-to-1 parity oblivious multiplexing
Robustness-based quantifiers are advantageous in general since there is a fair knowledge about their properties, for instance, that in many resource theories, they are monotonic under linear combinations of free operations [53]. Moreover, robustness has a strong operational appeal since it relates to the ubiquitous characteristic in laboratories of uncontrolled degrees of freedom, and hence it would be interesting if statements about robustness of contextuality to other types of noise could be produced for the parity-oblivious tasks. Dephasing noise, for instance, is universally present in quantum computation and represents a considerable obstacle to the implementation of many protocols, especially concerning scalability. Moreover, it displays similarities with depolarising noise, in the sense that complete dephasing will deem any prepare-and-measure scenario noncontextual [41]. In this section, we thus analyze robustness to dephasing and how this contextuality measure relates to quantifying advantage in POM tasks. However, general results such as Prop. 1 are challenging when it comes to dephasing since it is a more structured type of noise, making arguments of this type difficult to build. We will therefore rely on the numerical investigation of the proposed 3-to-1 POM scenarios to investigate the usefulness of robustness to dephasing in quantifying the nonclassical advantage for these tasks by employing the modified code (see Appendix B).
@@
We further show that robustness to dephasing can still perform as a good quantifier by exploiting the symmetry of this scenario and minimizing the robustness of contextuality to dephasing over all relevant dephasing bases. In particular, we here consider all available measurement basis $\hat{X},\hat{Y},\hat{Z}$. To gain intuition on why, notice that these axes will maximise the action of the noise over the volume of the set of states, requiring a smaller amount of noise to produce a greater change. Allowing for axes that align with a particular preparation, for instance, would possibly yield a larger value of robustness since the polyhedron from Fig. 2 would keep one of its diagonals unchanged, and the value of robustness with respect to this axis would be discarded by the minimisation. Indeed, we show the numerical investigation leading to this choice in Appendix C.2.
@@ Acknowledgements
We thank Bárbara Amaral for all the support and guidance on this manuscript, and Rafael Wagner, Giulio Camillo, and Anubhav Chaturvedi for the fruitful discussion. A.M.F. was supported by the São Paulo Research Foundation (FAPESP) via Project Numbers 2022/05455-5 (FAS-Masters) and 2022/13208-8 (RIA-Research Internship Abroad), and from the Instituto Serrapilheira, Grant 2021/4. V.P.R. was supported by the Foundation for Polish Science (IRAP project, ICTQT, contract no. MAB/2018/5, co-financed by EU within Smart Growth Operational Programme). J.H.S. was supported by the National Science Centre, Poland (Opus project, Categorical Foundations of the Non-Classicality of Nature, Project No. 2021/41/B/ST2/03149). R.D.B. and A.B.S. acknowledge support by the Digital Horizon Europe project FoQaCiA, Foundations of quantum computational advantage, GA No. 101070558, funded by the European Union, NSERC (Canada), and UKRI (UK). Some figures were prepared using TikZit.
@@ Appendix A Linear program for testing nonclassicality – Formal definitions
Alternatively, these states and effects can be naturally described with respect to the subspaces of $V$ they span (which will most probably not match [^6] Consider, for instance, the POM scenario introduced in Sec. I.3 for $\theta=\frac{\pi}{2}$. In this case, all states lie in the equator of the Bloch sphere, spanning $\mathbbm{R}^{2}$, while the effects remain lying on the three Pauli axes and therefore spanning $\mathbbm{R}^{3}$. This is only one of many examples in which the sets of states and effects in a fragment span different subspaces.), respectively $\mathsf{Span}(\Omega)$ and $\mathsf{Span}(\mathcal{E})$, and from here on we will always assume this is the case unless stated otherwise. One can easily describe these vectors with respect to the full vector space $V$ by employing inclusion maps $I_{\Omega}:\mathsf{Span}(\Omega)\to V$ and $I_{\mathcal{E}}:\mathsf{Span}(\mathcal{E})\to V$, in which case the tuple $(\Omega,\mathcal{E},I_{\Omega},I_{\mathcal{E}})$ is referred to as the accessible GPT fragment [62].
@@ Appendix B Modification of the code
The implementation for the linear program in Ref. [39] requests as inputs the sets of states and effects $(\Omega,\mathcal{E})$ that constitute the fragment, a privileged effect $u$ that plays the role of the discard effect and a privileged state $\mu$, called maximally mixed state. The code then constructs the accessible GPT fragment associated with the input fragment and the fully depolarising map $D_{depol}$, consisting of discarding any state it receives and preparing the maximally mixed state instead. The code in Mathematica has the additional feature of checking whether the input states and effects are represented by density operators and POVM element matrices, and if they do, already assumes the discard effect to be $\mathbbm{1}_{\mathcal{H}}$ and the maximally mixed state to be $\mathbbm{1}_{\mathcal{H}}/\text{dim}(\mathcal{H})$, where $\mathcal{H}$ is the Hilbert space with respect to which the density operators and POVM elements are represented. [^8] In this case, the user does not need to provide the discard and the maximally mixed state. Both implementations then proceed to characterise the accessible GPT fragment and its cone facets and solve Program 1. Finally, they construct an ontological model for the depolarised scenario by factorising the matrix $\sigma$ found by the program, and output an array $(r,\vec{\mu},\vec{\xi})$, where $r$ is the robustness of contextuality to depolarisation, $\vec{\mu}$ is a list of epistemic states for the noncontextual ontological model, and $\vec{\xi}$ a list of the respective response functions.
@@
Finally, as an example of our reasoning, consider a scenario that takes states and effects from a qubit and in which dephasing happens with respect to the $Z$ axis. The code will receive as input the set =$\mathcal{M}=\{\frac{\mathbbm{1}+\hat{Z}}{\sqrt{2}},\frac{\mathbbm{1}-\hat{Z}}{\sqrt{2}}\}$, and the dephasing map takes the form
@@
$D_{deph}\cdot\frac{1}{\sqrt{2}}\left(\begin{array}[]{c}\braket{\mathbbm{1}}\\ \braket{\hat{X}}\\ \braket{\hat{Y}}\\ \braket{\hat{Z}}\end{array}\right)=\frac{1}{\sqrt{2}}\left(\begin{array}[]{c}\braket{\mathbbm{1}}\\ 0\\ 0\\ \braket{\hat{Z}}\end{array}\right),$ (26)

### v3

@@ II.1 Robustness to depolarising noise in $n$-to-1 parity-oblivious multiplexing
$p(k=x_{y}|P_{x},M_{y})$ $=$ $\operatorname{Tr}(E_{k=x_{y}|M_{y}}\rho_{x})$ (15)
$\mapsto$ $(1-r)\operatorname{Tr}(E_{k=x_{y}|M_{y}}\rho_{x})$
$\qquad\qquad\quad+r\operatorname{Tr}(E_{k=x_{y}|M_{y}}\mu)$
$=$ $(1-r)p(k=x_{y}|P_{x},M_{y})$
$\qquad\qquad\quad+r\operatorname{Tr}(E_{k=x_{y}|M_{y}}\mu),$
where $E_{k=x_{y}|M_{y}}$ are the POVM elements of measurement $M_{y}$. Once we calculate the success rate for the depolarised version of the experiment, we find
$s_{depol}$ $=\frac{(1-r)}{2^{n}n}\sum_{y=1}^{n}\sum_{x\in[0,1]^{\times n}}p(k=x_{y}|P_{x},M_{y})$
$\qquad\qquad+\frac{r}{2^{n}n}\sum_{y=1}^{n}\sum_{x\in X_{n}}\operatorname{Tr}(E_{k=x_{y}|M_{y}}\mu)$ (16)
@@
where the second equality follows from the fact that for every choice of measurement $y$, there will be $2^{n-1}$ strings $x$ with $y$-th input equal to 00, and $2^{n-1}$ strings with the same input equal to $1$. Since all $E_{k=x_{y}|M_{y}}$ are elements of POVM, it must be that $E_{k=0_{y}|M_{y}}+E_{k=1_{y}|M_{y}}=\mathbb{1}$ for any $M_{y}$, simplifying the summation.
@@
With this relation, one can derive bounds on the number of bits that can be optimally encoded in a particular quantum system by assuming bounds on the maximal robustness to depolarising achievable for this system. For instance, for qubits, it is common to assume that $r_{min}^{depol}$ can never exceed the value $\frac{1}{2}$ for whatever signature of quantum advantage taken into consideration, in particular $r_{min}^{depol}<\frac{1}{2}$ for a finite amount of extremal states and effects in the fragment (which is always the case in $n$-to11-1- 1 POM tasks with finite $n$). We provide a compelling numerical induction for this bound in C.1 that relies on the simplex embeddability of an ever-increasing set of preparations and measurements. Then notice that the optimal quantum success rate is given by $s=\frac{1}{2}(1+\frac{1}{\sqrt{n}})$ [13], while $s_{NC}=\frac{1}{2}(1+\frac{1}{n})$ [4]. If $r_{min}^{depol}<\frac{1}{2}$, then
@@
This result has been derived before by geometrical [47] and locality [48] arguments, but Proposition 1 allows us to reframe it solely on simplicial embedding assumptions. This method can, in principle, be generalised: if one finds bounds on robustness for quantum systems of higher dimensions, one can immediately use Eq. (19) to derive an upper bound to the number of optimally encoded bits.
@@ II.2 Robustness to dephasing noise in 3-to-1 parity oblivious multiplexing
Robustness-based quantifiers are advantageous in general since there is a fair knowledge about their properties, for instance, that in many resource theories, they are monotonic under linear combinations of free operations [53]. Moreover, robustness has a strong operational appeal since it relates to the ubiquitous characteristic in laboratories of uncontrolled degrees of freedom, and hence it would be interesting if statements about robustness of contextuality to other types of noise could be produced for the parity-oblivious tasks. Dephasing noise, for instance, is universally present in quantum computation and represents a considerable obstacle to the implementation of many protocols, especially concerning scalability. Moreover, it displays similarities with depolarising noise, in the sense that complete dephasing will deem any prepare-and-measure scenario noncontextual [40]. In this section, we thus analyze robustness to dephasing and how this contextuality measure relates to quantifying advantage in POM tasks. However, general results such as Prop. 1 are challenging when it comes to dephasing since it is a more structured type of noise, making arguments of this type difficult to build. We will therefore rely on the numerical investigation of the proposed 3-to-1 POM scenarios to investigate the usefulness of robustness to dephasing in quantifying the nonclassical advantage for these tasks by employing the modified code (see B).
@@
We further show that robustness to dephasing can still perform as a good quantifier by exploiting the symmetry of this scenario and minimizing the robustness of contextuality to dephasing over all relevant dephasing bases. In particular, we here consider all available measurement basis $\hat{X},\hat{Y},\hat{Z}$. To gain intuition on why, notice that these axes will maximise the action of the noise over the volume of the set of states, requiring a smaller amount of noise to produce a greater change. Allowing for axes that align with a particular preparation, for instance, would possibly yield a larger value of robustness since the polyhedron from Fig. 2 would keep one of its diagonals unchanged, and the value of robustness with respect to this axis would be discarded by the minimisation. Indeed, we show the numerical investigation leading to this choice in C.2.
@@ Acknowledgements
We thank Bárbara Amaral for all the support and guidance on this manuscript, and Rafael Wagner, Giulio Camillo, and Anubhav Chaturvedi for the fruitful discussion. A.M.F. was supported by the São Paulo Research Foundation (FAPESP) via Project Numbers 2022/05455-5 (FAS-Masters) and 2022/13208-8 (RIA-Research Internship Abroad), and from the Instituto Serrapilheira, Grant 2021/4. V.P.R. was supported by the Foundation for Polish Science (IRAP project, ICTQT, contract no. MAB/2018/5, co-financed by EU within Smart Growth Operational Programme). J.H.S. was supported by the National Science Centre, Poland (Opus project, Categorical Foundations of the Non-Classicality of Nature, Project No. 2021/41/B/ST2/03149). V.P.R., R.D.B., and A.B.S. acknowledge support by the Digital Horizon Europe project FoQaCiA, Foundations of quantum computational advantage, GA No. 101070558, funded by the European Union, NSERC (Canada), and UKRI (UK). Some figures were prepared using TikZit and Mathcha. This research was partially conducted while visiting the Okinawa Institute of Science and Technology (OIST) through the Theoretical Sciences Visiting Program (TSVP).
@@ Appendix A Linear program for testing nonclassicality – Formal definitions
Alternatively, these states and effects can be naturally described with respect to the subspaces of $V$ they span (which will most probably not match [^7] Consider, for instance, the POM scenario introduced in Sec. I.3 for $\theta=\frac{\pi}{2}$. In this case, all states lie in the equator of the Bloch sphere, spanning $\mathbb{R}^{2}$, while the effects remain lying on the three Pauli axes and therefore spanning $\mathbb{R}^{3}$. This is only one of many examples in which the sets of states and effects in a fragment span different subspaces.), respectively $\mathsf{Span}(\Omega)$ and $\mathsf{Span}(\mathcal{E})$, and from here on we will always assume this is the case unless stated otherwise. One can easily describe these vectors with respect to the full vector space $V$ by employing inclusion maps $I_{\Omega}:\mathsf{Span}(\Omega)\to V$ and $I_{\mathcal{E}}:\mathsf{Span}(\mathcal{E})\to V$, in which case the tuple $(\Omega,\mathcal{E},I_{\Omega},I_{\mathcal{E}})$ is referred to as the accessible GPT fragment [62].
@@ Appendix B Modification of the code
The implementation for the linear program in Ref. [38] requests as inputs the sets of states and effects $(\Omega,\mathcal{E})$ that constitute the fragment, a privileged effect $u$ that plays the role of the discard effect and a privileged state $\mu$, called maximally mixed state. The code then constructs the accessible GPT fragment associated with the input fragment and the fully depolarising map $D_{depol}$, consisting of discarding any state it receives and preparing the maximally mixed state instead. The code in Mathematica has the additional feature of checking whether the input states and effects are represented by density operators and POVM element matrices, and if they do, already assumes the discard effect to be $\mathbb{1}_{\mathcal{H}}$ and the maximally mixed state to be $\mathbb{1}_{\mathcal{H}}/\text{dim}(\mathcal{H})$, where $\mathcal{H}$ is the Hilbert space with respect to which the density operators and POVM elements are represented. [^9] In this case, the user does not need to provide the discard and the maximally mixed state. Both implementations then proceed to characterise the accessible GPT fragment and its cone facets and solve Program 1. Finally, they construct an ontological model for the depolarised scenario by factorising the matrix $\sigma$ found by the program, and output an array $(r,\vec{\mu},\vec{\xi})$, where $r$ is the robustness of contextuality to depolarisation, $\vec{\mu}$ is a list of epistemic states for the noncontextual ontological model, and $\vec{\xi}$ a list of the respective response functions.
@@
Finally, as an example of our reasoning, consider a scenario that takes states and effects from a qubit and in which dephasing happens with respect to the $Z$ axis. The code will receive as input the set =$\mathcal{M}=\{\frac{\mathbb{1}+\hat{Z}}{\sqrt{2}},\frac{\mathbb{1}-\hat{Z}}{\sqrt{2}}\}$, and the dephasing map takes the form
@@
$D_{deph}\cdot\frac{1}{\sqrt{2}}\left(\begin{array}[]{c}\braket{\mathbb{1}}\\ \braket{\hat{X}}\\ \braket{\hat{Y}}\\ \braket{\hat{Z}}\end{array}\right)=\frac{1}{\sqrt{2}}\left(\begin{array}[]{c}\braket{\mathbb{1}}\\ 0\\ 0\\ \braket{\hat{Z}}\end{array}\right),$ (34)